OpenStudy (anonymous):

16= 1/2 4(3^1+b^2) HELP HELP HELP NOW PLEASE??? the ^ are supposed to be aiming down so that the exponents are not up.

OpenStudy (anonymous):

Step 1: Simplify both sides of the equation. 16=2b2+6 Step 2: Subtract 2b^2 from both sides. 16−2b2=2b2+6−2b2 −2b2+16=6 Step 3: Subtract 16 from both sides. −2b2+16−16=6−16 −2b2=−10 Step 4: Divide both sides by -2. −2b2 −2 = −10 −2 b2=5 Step 5: Take square root. b=±√5 b=√5 or b=−√5 Answer: b=√5 or b=−√5

OpenStudy (anonymous):

Thank you!